Maya was brother was 12 when she was born. The sum of their ages is 22. Find their ages.
Answer by Maths68(1474)

Maya was brother was 12 when she was born. The sum of their ages is 22. Find their ages.
Let
Present age of Maya = m
Present age of Brother = b
Maya is 12 years older than her brother
m=b+12.....(1)

Sum of their ages is 22
m+b=22.....(2)
Put the value of m from (1) to (2)
b+12+b=22
2b=22-12
2b=10
b=10/2
b=5
Plug in the value of b in (1)
m=5+12
m=17

Present age of Maya = m = 17 years old
Present age of Brother = b = 5 years old